# What Should We Engineer in Prompts? Training Humans in Requirement-Driven LLM Use

### v2

@@ 6.2. In-depth Analysis: The Validity of ROPE Paradigm
Table 3. Spearman correlations ($\rho$) between LLM Output Quality and Requirement Quality scores by tasks
Task Type Task $\rho$
GPTs Task TripAdvisor 0.79
OutlineAssistant 0.75
Game Task TicTacToe 0.90
Connect4 0.08
Overall 0.71
Going beyond learning gains, we dive deeper into the connections between requirements in prompts and LLM output quality, as well as the role of optimizers.
@@
To examine whether more correct and complete requirements led to better LLM outputs, we calculated the correlation between the two components of our Overall Scores: Requirement Quality and LLM Output Quality scores. We found a strong positive correlation, with a Spearman’s correlation coefficient of $\rho=0.71$ (Schober et al., 2018). However, task-specific analysis revealed nuances. As shown in Table 3, while most tasks achieved $\rho\geq 0.7$, Connect4 showed no correlation. Upon further inspection, we found that interestingly, the LLM tends to produce two implementation types for Connect4: command line and pygame. Prompts that include interactive or visual requirements such as ‘‘Highlight the hovered column’’ often led to pygame implementations. If we split the Connect4 correlations by the LLM output implementation type, we get reasonable numbers for both cases: $\rho=0.89$ (command line) and $\rho=0.58$ (pygame), suggesting that implementations may respond to requirements differently. For example, the requirement ‘‘Display player’s turn message’’ was always implemented in command line Connect4, even if omitted in prompts. However, only 1 pygame Connect4 implemented the same requirement, out of 6 participants who correctly described it. This suggests that requirements can be interdependent, with LLM-hardness varying by implementations.
We further analyzed how different types of requirement defects impact LLM output quality. We found that the number of omission errors (incomplete requirements) had a stronger negative impact on LLM Output Quality Score ($\rho=-0.49$) compared to commission errors (inaccurate requirements) ($\rho=0.03$). This suggests that while LLMs can correct inaccuracies in requirements, they struggle to fill in missing information — further highlighting the importance of training humans to express all their necessary requirements.
@@
Finally, we evaluated how the optimizer affected prompt quality, experimenting on a single optimizer to explore the feasibility, without generalizing to all optimizers (Section 7.1).
@@
Importantly, the optimizer could not close the performance gap between the ROPE and PE group. We computed the gains of PE participants after using the optimizer ($7.3\%$ = optimized prompts’ post-test scores -- original prompts’ pre-test scores). Using a two-sample t-test, the ROPE’s 19.0% pre- to post-test gains remained significantly higher ($p=0.01$) than PE group’s optimized gains ($7.3\%$). This supports the finding that missing, highly customized requirements are unlikely to be added automatically.
While the optimizer generally had a positive impact, we also observed instances where it misinterpreted user inputs or added irrelevant details. For instance, for four out of five participants who requested an alphabetic list (A, B, C), the optimizer changed it to a numbered list. In some cases, these misinterpretations drastically reduced LLM performance. For example, two participants used the word “interactive” in their prompt, leading the optimizer to incorrectly add the role of “intensive interactive web application developer” and request the use of “Flask and HTML” in the optimized prompts, leading to poorly implemented code since it is much harder to generate a functional web app.
These issues highlight potential gaps in user-expressed requirements and LLM-interpreted requirements, suggesting that future training should address the risks of misinterpretation by either the LLM or the optimizer (Section 7.2). On a more positive note, such misinterpretations could provide useful feedback in the future, alerting users to potential misalignment between their requirements and the LLM’s understanding.
@@ 7. Discussion
In this work, we introduce Requirement-Oriented Prompt Engineering (ROPE), a human-AI collaborative prompting paradigm where humans focus on effective requirements specification. Our evaluation shows that requirement-focused training significantly enhances novices’ ability to extract and articulate requirements in prompts, leading to more goal-aligned LLM outputs; we also notice a mismatch between humans’ under-specified requirements and LLMs’ misinterpretations is a key communication barrier (Section 6.2). While our work made an important step toward ROPE, several limitations remain, and there are many interesting questions to explore. Here, we reflect on possible immediate extensions on requirement-focused training, as well as the longer-term evolution of the ROPE paradigm.
@@ 7.1. Limitations and Next Steps on Requirement-Focused Training
In addition to extending length, our training can also be broadened for better generalizability. Our study demonstrated that the requirement specification skills users acquired through training generalized to new GPTs and Game development tasks. However, an open question remains: how broadly do these requirement skills transfer to other natural language programming tasks (e.g., data analysis or creative content generation)? In preparation for generalizability, our ROPE interface is adaptable to different tasks, as the training and feedback loop can be easily generated based on provided reference requirements and LLM outputs. We open-source ROPE system and encourage its use for customized tasks across domains. Moreover, beyond human-AI interactions, ROPE skills — such as requirement extraction and iterative refinement — can be broadly applicable to fields like software engineering and design. Future research should explore further transfer tasks in assessments, and also how ROPE skills may enhance communication, problem decomposition, and computational thinking across various fields.
Another dimension of generalizability is towards users’ different prompting behaviors. Our study focused on extracting fixed requirements, which allowed for controlled feedback and simulated the scenarios when interaction examples are available. However, this setup did not fully capture all real-world prompting scenarios: users often engage in one-off interactions (e.g., direct question-answering) or iterative and open-ended prompting (e.g., chatting with LLMs on a rough idea), etc. (Shankar et al., 2024a; Kim et al., 2024). Future studies should investigate more self-directed tasks, compare different types of prompting, and tailor training accordingly. While providing feedback without clear requirements may be challenging, pre-generated materials on common novice approaches could offer useful scaffolding.
One primary bottleneck for broadening the scope is our lack of automatic assessment methods for the LLM-hard tasks. Our time-consuming manual grading effort limited our ability to conduct more experiments by e.g., grading multiple LLM outputs using different temperatures per user prompt, experimenting with multiple optimizers, etc. Future work should explore automatic assessments and their trade-offs. For example, LLM-as-a judge paired with few-shot prompting might provide reasonable estimations (Zhao et al., 2021) and might be suitable for tasks that contain substantial numbers of requirements where wrong grading on a single requirement does not drastically affect the assessment results.
@@ 7.2. Reflection on the ROPE Paradigm
Along with prior work on requirement-oriented evaluation (Kim et al., 2024), we believe that ROPE moves us towards a future where requirements serve as the central interface between AI and humans. Here, we discuss key skills humans and LLMs need to develop to prepare for a ROPE future. These insights can also be useful for designing better prompting support tools.
@@
Additionally, users should iterate requirement granularity based on task “LLM-hardness.” This involves observing LLM failures and refining requirements, essentially developing a theory of mind for LLMs (Wang et al., 2024b). Some of our participants were already thinking about “what is obvious enough for ChatGPT to already understand/perform well on” (S28) and adjust their prompt accordingly (although not necessarily correctly). In our training, we introduced the concept of requirement specificity in Tetris, where users generate requirements in two levels of granularity (i.e., main steps and details, Section 4.3). However, some tasks required even greater specificity, such as generating non-standard Connect4 gameplay elements ‘‘Cross out winning cells’’. This ability to adjust specificity in requirements can be influenced by domain expertise, for example, expert developers can iterate prompts with more accurate details to successfully generate code while students often struggle (Nam et al., 2023). Future studies should offer adaptive support for adjusting requirement granularity based on task complexity and user expertise.
@@
Second, optimizers can also assist test data synthesis. More advanced optimizers need to iterate prompts on validation datasets and automatic objective functions, similar to machine learning. Insights from software engineering test automation or NLP approaches to extract or synthesize test cases from natural language requirements could be helpful (Garousi et al., 2020; Yang et al., 2023; Yu et al., 2024; Yang et al., 2024b).
Finally, LLMs or optimizers should explicitly communicate their interpretations and reveal biases in their implementation. As we observed in our study, optimizers can misinterpret requirements or add unnecessary details (e.g., “interactive game” can lead to unwanted features like Flask integration). LLMs should explicitly communicate their understandings of user intent, such as translating ambiguous natural language requirements into more formal specifications (Liu et al., 2023b; Lee et al., 2022), asking clarifying questions (Qian et al., 2024; Lahiri et al., 2022; Chen et al., 2024), or extracting and refining generated criteria (Kim et al., 2024; Yuan et al., 2024; Wang et al., 2024a). This is not a perfect solution, as models’ translations may still reflect their biases and models may miss to ask about LLM-hard requirements. Nonetheless, to correctly implement requirements in LLM outputs when what is LLM-hard keeps evolving, humans need to learn when and how to specify requirements with more granularity, and models need to keep communicating what requirements might be LLM-hard or LLM-conflict.
@@ Acknowledgements.
Thanks to all the participants for the pilot, interview, and user study in this work. Thanks to Kelly Rivers, Michael Taylor, Michael Hilton, Michael Xieyang Liu, Xinran Zhao, Lauren Sands, Austin Schick, David Kosbie, and Daniel Anderson for all the insights, advice, and help. Thanks to the CMU HCII faculty and Ken’s lab for feedback. Thanks to the National Science Foundation (award CNS-2213791, 2414915) and gift fund from Google for support of this work. Thanks to the OpenAI research credit program and Amazon AI research gift fund.

### v3

@@ 6.2. In-depth Analysis: The Validity of ROPE Paradigm
Going beyond learning gains, we dive deeper into the connections between requirements in prompts and LLM output quality, the role of optimizers, as well as the differences brought by model advancement.
@@
To examine whether more correct and complete requirements led to better LLM outputs, we calculated the correlation between the two components of our Overall Scores: Requirement Quality and LLM Output Quality scores using two LLMs generating the outputs — GPT-4o as in our main experiment, and o3-mini as explained below. We found a strong positive correlation, with a Spearman’s correlation coefficient of $\rho=0.71$ (Schober et al., 2018). However, task-specific analysis revealed nuances. As shown in Table 3, while most tasks achieved $\rho\geq 0.7$, Connect4 showed no correlation when LLM output is generated using GPT-4o. Upon further inspection, we found that interestingly, GPT-4o tends to produce two implementation types for Connect4: command line and pygame. Prompts that include interactive or visual requirements such as ‘‘Highlight the hovered column’’ often led to pygame implementations. If we split the Connect4 correlations by the LLM output implementation type, we get reasonable numbers for both cases: $\rho=0.89$ (command line) and $\rho=0.58$ (pygame), suggesting that implementations may respond to requirements differently. For example, the requirement ‘‘Display player’s turn message’’ was always implemented in command line Connect4, even if omitted in prompts. However, only 1 pygame Connect4 implemented the same requirement, out of 6 participants who correctly described it. This suggests that requirements can be interdependent, with LLM-hardness varying by implementations.
Table 3. Spearman correlations ($\rho$) between LLM Output Quality and Requirement Quality scores by tasks and models.
Task $\rho$ (GPT-4o) $\rho$ (o3-mini)
GPTs TripAdvisor 0.79 0.75
OutlineAssistant 0.75 0.83
Game TicTacToe 0.90 0.90
Connect4 0.10 0.67
Overall 0.71 0.80
We further analyzed how different types of requirement defects impact LLM output quality. We found that the number of omission errors (incomplete requirements) had a stronger negative impact on LLM Output Quality Score ($\rho=-0.49$) compared to commission errors (inaccurate requirements) ($\rho=0.05$). This suggests that while LLMs can correct inaccuracies in requirements, they struggle to fill in missing information — further highlighting the importance of training humans to express all their necessary requirements.
@@
To assess how the optimizer affected prompt quality, we experimented on a single optimizer to explore the feasibility, without generalizing to all optimizers (Section 7.1).
@@
Importantly, the optimizer could not close the performance gap between the ROPE and PE group. We computed the gains of PE participants after using the optimizer ($7.3\%$ = optimized prompts’ post-test scores -- original prompts’ pre-test scores). Using a two-sample t-test, the ROPE’s 19.1% pre- to post-test gains remained significantly higher ($p=0.01$) than PE group’s optimized gains ($7.3\%$), as shown in Figure 6. This supports the finding that missing, highly customized requirements are unlikely to be added automatically.
Figure 6. The pre-test and post-test overall scores between PE and ROPE conditions and the original and optimized prompt versions. ROPE’s learning gain remains significantly higher than PE’s optimized gains (** denotes $p\leq 0.01$).
[…]
Bar chart comparing pre-test and post-test scores for original and optimized prompts in both PE and ROPE conditions. ROPE’s original learning gains still outperform PE’s optimized gains significantly (p = 0.01).
While the optimizer generally had a positive impact, it might misinterpret user inputs or add irrelevant details. For instance, for four out of five participants who requested an alphabetic list (A, B, C), the optimizer changed it to a numbered list. In some cases, these misinterpretations drastically reduced LLM performance. For example, two participants used the word “interactive” in their prompt, leading the optimizer to incorrectly add the role of “intensive interactive web application developer” and request the use of “Flask and HTML” in the optimized prompts, leading to poorly implemented code since it is much harder to generate a functional web app.
These issues highlight potential gaps in user-expressed requirements and LLM-interpreted requirements, suggesting that future training should address the risks of misinterpretation by either LLMs or optimizers (Section 7.2). On a more positive note, such misinterpretations could provide useful feedback, alerting users to potential misalignment between their requirements and the LLM’s understanding.
ROPE’s effectiveness generalizes to more advanced (reasoning) LLMs.
Noticing the continuous evolvement of LLMs, we naturally ask the question: Would ROPE’s effectiveness generalize beyond the GPT-4o we experimented, to models with better capabilities and alignments to human intents? Recent advancements in LLMs have introduced reasoning models, which have improved instruction-following capabilities and exhibit better performance in complex reasoning tasks (DeepSeek-AI et al., 2025). We conducted an additional experiment using OpenAI’s o3-mini, a representative model of the emerging reasoning LLMs that has been shown to surpass GPT-4o (Chiang et al., 2024). We had o3-mini generate outputs using the same prompts (both original and optimized ones as above), and repeated all the grading and analyses as done on GPT-4o.
The results showed that ROPE-trained prompts not only remained effective, but in many cases performed even better with o3-mini. Specifically, when using the overall scores calculated on o3-mini (which impacted the LLM Output Quality score), we saw a larger learning gains compared to GPT-4o (23% $>$ 19% for the ROPE group, 2% $>$ 1% for PE group). This learning gap again could not be closed by the optimizer. Digging deeper, we found that o3-mini improved the LLM Output Quality score by 9.7% compared to GPT-4o (paired t-test, $p<0.001$), which means that o3-mini better implemented the prompts. This suggests that improved model capabilities do not render explicit human requirements obsolete, but rather amplify their benefits, making learning on requirements more reflective on the model output quality.
Furthermore, we also found a stronger alignment between user-specified requirements and generated outputs for o3-mini, indicated by the stronger correlation between Requirement Quality and LLM Output Quality scores when using o3-mini compared to GPT-4o ($\rho=0.80>0.71$, as shown in Table 3). The largest improvement in correlation came from the Connect4 task, where o3-mini showed better abilities to implement customized requirements like ‘‘Highlight the hovered column.’’ These findings underscore the continued relevance of human-authored requirements even as LLMs evolve (more discussions in Section 7.2).
@@ 7. Discussion
In this work, we introduce Requirement-Oriented Prompt Engineering (ROPE), a human-AI collaborative prompting paradigm where humans focus on effective requirements specification. Our evaluation shows that requirement-focused training significantly enhances novices’ ability to extract and articulate requirements in prompts, leading to more goal-aligned LLM outputs; we also notice a key communication barrier between humans’ under-specified requirements and LLMs’ misinterpretations (Section 6.2).
While our work made an important step toward ROPE, several limitations remain, and there are many interesting questions to explore. Here, we reflect on possible immediate extensions on requirement-focused training, as well as the longer-term evolution of the ROPE paradigm.
@@ 7.1. Limitations and Next Steps on Requirement-Focused Training
In addition to extending length, our training can also be broadened for better generalizability. Our study demonstrated that the requirement specification skills users acquired through training generalized to new GPTs and Game development tasks. However, an open question remains: how broadly do these requirement skills transfer to other natural language programming tasks (e.g., data analysis or creative content generation)? Does training on easier tasks generalize to harder tasks? In preparation for generalizability, our ROPE interface is adaptable to different tasks, as the training and feedback loop can be easily generated based on provided reference requirements and LLM outputs. We open-source ROPE system and encourage its use for customized tasks across domains. Moreover, beyond human-AI interactions, ROPE skills — such as requirement extraction and iterative refinement — can be broadly applicable to fields like software engineering and design. Future research should explore further transfer tasks in assessments, and also how ROPE skills may enhance communication, problem decomposition, and computational thinking across various fields.
Another dimension of generalizability is towards users’ different prompting behaviors. Our study focused on extracting fixed requirements, which allowed for controlled feedback and simulated the scenarios when interaction examples are available. However, this setup did not fully capture all real-world prompting scenarios: users often engage in one-off interactions (e.g., direct question-answering) or iterative and open-ended prompting (e.g., chatting with LLMs on a rough idea) (Shankar et al., 2024a; Kim et al., 2024). Future studies should investigate more self-directed tasks, compare different types of prompting, and tailor training or support accordingly. While providing feedback without clear requirements may be challenging, pre-generated materials on common novice approaches could offer useful scaffolding.
One primary bottleneck for broadening the scope is our lack of automatic assessment methods for the LLM-hard tasks. Our time-consuming manual grading effort limited our ability to conduct more experiments by e.g., grading multiple LLM outputs using different temperatures per user prompt, experimenting with multiple optimizers, etc. Future work should explore automatic assessments and their trade-offs. For example, LLM-as-a judge paired with few-shot prompting might provide reasonable estimations (Zhao et al., 2022) and might be suitable for tasks that contain substantial numbers of requirements where wrong grading on a single requirement does not drastically affect the assessment results. Future work can also explore alternative instructional designs such as using human experts or incorporate additional learning principles (Koedinger et al., 2013) to maximize the efficacy of ROPE training.
@@ 7.2. Reflection on the ROPE Paradigm
Along with prior work on requirement-oriented evaluation (Kim et al., 2024), we believe that ROPE moves us towards a future where requirements serve as the central interface between AI and humans especially as models become more aligned with humans — as we have seen in Section 6.2, more advanced models better implement user requirements. Here, we further discuss key skills humans and LLMs need to develop to prepare for a ROPE future. While our paper focused on a ROPE training system, our insights can also be useful for designing better prompting support tools.
@@
Additionally, users should iterate requirement granularity based on task “LLM-hardness.” This involves observing LLM failures and refining requirements, essentially developing a theory of mind for LLMs (Wang et al., 2024b). Some of our participants were already thinking about “what is obvious enough for ChatGPT to already understand/perform well on” (S28) and adjust their prompt accordingly (although not necessarily correctly). In our training, we introduced the concept of requirement specificity in Tetris, where users generate requirements in two levels of granularity (i.e., main steps and details, Section 4.3). However, some tasks required even greater specificity, such as generating non-standard Connect4 gameplay elements ‘‘Cross out winning cells’’. This ability to adjust specificity in requirements can be influenced by domain expertise, for example, expert developers can iterate prompts with more accurate details to successfully generate code while students often struggle (Nam et al., 2024). Future studies should offer adaptive support for adjusting requirement granularity based on task complexity and user expertise, such as via interactive dialog and automated diagnosis.
We believe that improving humans’ requirement specification skills will remain important despite rapid model advancements, since it is grounded in a fundamental challenge in human–LLM interaction: LLMs need access to user intent, but specific intent may not be available until users articulate it. As mentioned before, requirements might be ambiguous or entirely missing. The former might be mitigated by more capable models (or prompting support tools) that are better at e.g., making personalized guesses on user needs, asking clarification questions, or providing requirement-specific feedback (Mu et al., 2024; Zhang and Choi, 2023). However, these can only be complementary to humans’ own requirement elicitation — a model that always asks for clarification on vague instructions can quickly become tedious and overwhelming to interact with. More importantly, missing requirements will remain a challenge, especially in LLM-hard tasks. When humans actually need to deviate the model’s behavior but forget to articulate it, models may already have a strong default behavior that prevents them from noticing anything missing. As a result, missing requirements risk becoming unknown unknowns (Lakkaraju et al., 2017; Rumsfeld, 2011) to LLMs, and models may default to “common sense” completions that conflict with users’ actual needs (e.g., defaulting to circles instead of X and O’s in customized Connect4 games). Given decades of research in design and software engineering show that requirement elicitation is non-trivial for humans (Section 2.3), we anticipate ROPE training will remain essential, even when LLMs and optimizers become more powerful.
@@
Second, optimizers can also assist test data synthesis. More advanced optimizers need to iterate prompts on validation datasets and automatic objective functions, similar to machine learning. Insights from software engineering test automation or NLP approaches to extract or synthesize test cases from natural language requirements could be helpful (Garousi et al., 2020; Yang et al., 2023; Yu et al., 2023; Yang et al., 2024a).
Finally, LLMs or optimizers should explicitly communicate their interpretations and reveal biases in their implementation. As we observed in our study, optimizers can misinterpret requirements or add unnecessary details (e.g., “interactive game” can lead to unwanted features like Flask integration). LLMs should explicitly communicate their understandings of user intent, such as translating ambiguous natural language requirements into more formal specifications (Liu et al., 2023a; Lee et al., 2022), asking clarifying questions (Qian et al., 2024; Lahiri et al., 2022; Chen et al., 2024), or extracting and refining generated criteria (Kim et al., 2024; Yuan et al., 2024; Wang et al., 2024a). This is not a perfect solution, as models’ translations may still reflect their biases and models may miss to ask about LLM-hard requirements. Nonetheless, to correctly implement requirements in LLM outputs when what is LLM-hard keeps evolving, humans need to learn when and how to specify requirements with more granularity, and models need to keep communicating what requirements might be LLM-hard or LLM-conflict.
@@ Acknowledgements.
Thanks to all the participants for the pilot, interview, and user study in this work. Thanks to Kelly Rivers, Michael Taylor, Michael Hilton, Michael Xieyang Liu, Xinran Zhao, Lauren Sands, Austin Schick, David Kosbie, and Daniel Anderson for all the insights, advice, and help. Thanks to the CMU HCII faculty and Ken’s lab for feedback. Thanks to the National Science Foundation (award CNS-2213791, 2414915) and gift fund from Google for support of this work. Thanks to the OpenAI research credit program and Amazon AI research gift fund. Thanks to all reviewers of this work.